OpenStudy (anonymous):

find a formula for the inverse g(x)=(x-6)^3

OpenStudy (lgbasallote):

first change g(x) into y \[\implies y = (x-6)^3\] now interchange x and y \[\implies x = (y - 6)^3\] solve for y \[\implies \sqrt[3] x = y -6\] add 6 to both sides \[\implies \sqrt[3] x + 6 = y\] does that help?